OpenStudy (anonymous):

what is 22/125 as a decimal

OpenStudy (kymber):

Divide 22 by 125 :P

OpenStudy (accessdenied):

If I didn't have a calculator, I would have noted that the denominator is just 5*5*5. If we multiply each 5 by 2, making them all 10's, we'd have 1000 as the denominator. This is easier to deal with in division. We just multiply the numerator by 2*2*2 as well. (It's easier to do the multiplication by hand than division, to me) \[ \frac{22 \times 2*2*2}{125\times2*2*2} \\ = \frac{22 \times 8}{1000} \\ = \frac{176}{1000} \qquad \quad \text{division by 1000, move decimal to left 3 spots}\\ = 0.176 \] Of course, this wouldn't be as easy in all cases.